# Physics-informed neural networks for solving moving interface flow problems using the level set approach

### v1

@@ Abstract
This paper advances the use of physics-informed neural networks (PINNs) architectures to address moving interface problems via the level set method. Originally developed for other PDE-based problems, we particularly leverage PirateNet’s features—including causal training, sequence-to-sequence learning, random weight factorization, and Fourier feature embeddings—and tailor them to achieve superior performance in modeling interface dynamics. Numerical experiments validate this framework on benchmark problems such as Zalesak’s disk rotation and time-reversed vortex flow. We demonstrate that PINNs can efficiently solve level set problems exhibiting significant interface deformation without the need for upwind numerical stabilization, as generally required by classic discretization methods. Additionally, geometric reinitialization or mass conservation schemes have been revealed as unnecessary for accurate and efficient solutions. However, incorporating an Eikonal regularization term in the loss function with an appropriate weight can further enhance results in specific scenarios. Our results indicate that PINNs with the PirateNet architecture surpass conventional PINNs in accuracy, achieving state-of-the-art error rates of $L^{2}=0.14\%$ for Zalesak’s disk and $L^{2}=0.85\%$ for the time-reversed vortex flow problem, as compared to reference solutions.
@@ 1 Introduction
Our study begins by testing our PINN approach on Burgers’ equation in 1D and 2D [43]. We then address level set problems, starting with a fabricated simple uniform transport case to highlight the challenges faced by the standard PINN solver and demonstrate the advantages of the proposed techniques. Next, we tackle two well-known level set problems: the rigid-body rotation of Zalesak’s disk [44] and a more complex time-reversed vortex flow problem, which features a varying velocity field [10]. For each case, we compare the reference solutions to results obtained from (a) the original PINN formulation [11], (b) the improved PINN approach [38], and (c) the state-of-the-art PirateNet framework [39]. Our findings demonstrate that PirateNet can solve complex level set problems effectively without requiring reinitialization or mass conservation schemes, making it a robust and efficient method for addressing such challenges.
@@
    We extend recent advanced PINN architectures to solve the level set equation. The PINNs’ loss function is augmented by two terms related to the residuals of the Eikonal equation and the mass conservation constraint respectively.
@@
    With experiments on classical level set benchmarks, we show that PINNs can learn the level set equations without the need for upwind numerical stabilization, geometric reinitialization or mass conservation schemes.
@@ 3.2.1 PirateNet architecture
where each element of the matrix F is sampled from a Gaussian distribution $\mathcal{N}(0,\sigma^{2})$, and $\sigma$ is a hyperparameter. Random Fourier features are particularly effective in reducing spectral bias, enabling improved predictions for high-frequency components in the solution [27].
@@ 3.2.2 Random weight factorization
The implementation details of the RWF are as follows [38]. First, the neural network parameters are initialized using a standard scheme, such as the Glorot initialization [55]. Next, for each weight matrix $\bm{W}$, a scale vector $\bm{s}$ is sampled from a multivariate normal distribution $\mathcal{N}(\mu,\sigma I)$, with $\mu$ and $\sigma$ are hyperparameters, and $I$ denotes the identity matrix. At initialization, the weight matrix is factorized as $\bm{W}=\text{diag}(\exp(\bm{s}))\cdot\bm{V}$, where $\bm{V}$ represents the remaining factorized parameter. Finally, gradient descent optimization is applied directly to the new parameters, $\bm{s}$ and $\bm{V}$. According to [32], a suitable selection of RWF hyperparameters can shorten the distance between the initialization and the global minimum in the factorized parameter space relative to the original parameter space within their respective loss landscapes. Furthermore, [38] has demonstrated through numerical experiments that, in the case of PINNs, the RWF technique enhances both accuracy and training performance across a range of PDE benchmarks.
@@ 4 Results
In the following sections, we present the numerical results obtained from the proposed PINN solver for two well-known level set benchmarks. Before diving into these benchmarks, we validate our PirateNetsframework using two challenging benchmark problems in fluid dynamics, the 1D and 2D Burgers’ equations. Next, we focus on two level set benchmarks: Zalesak’s disk and the time-reversed vortex flow.
@@
4.1 Burgers’ 1D
$u(x=-L,t)$ $=u(x=L,t)=0,$
$u(x,t=0)$ $=0$
Configuration name Relative $L^{2}$-error (%)
PirateNet 0.070
No Causal Training 0.072
Default 0.073
No RWF 0.073
No Fourier Feature 0.093
No Grad Norm 0.151
Plain 0.496
4.2 Burgers’ 2D
$u(x,y,t=0)=\sin(2\pi x)\sin(2\pi y),$ (44)
$v(x,y,t=0)=\sin(\pi x)\sin(\pi y),$ (45)
$u(x,y,t)=0\quad\text{for}\quad(x,y)\in\partial\Omega,$ (46)
$v(x,y,t)=0\quad\text{for}\quad(x,y)\in\partial\Omega,$ (47)
      Plain       2.20       1.46       2.64       57
      Default       2.11       1.40       2.53       128
      No Fourier feature       2.12       1.40       2.54       100
      No RWF       2.10       1.40       2.53       108
      No grad norm       2.11       1.40       2.53       127
      No causal training       2.10       1.39       2.52       148
      No modified MLP       2.11       1.40       2.52       61
      PirateNet        2.10       1.39       2.52       168
Table 2: Burgers’ 2D: Ablation study.
(a) $u$ velocity fields at $t=[0,0.25,0.5]$s
(b) $v$ velocity fields at $t=[0,0.25,0.5]$s
It is also worth noting the impact of certain features on computational load. Although run times can vary due to the state of the hardware and computational nodes of the compute cluster, there is a significant increase in run time when comparing the Plain configuration with most other configurations with added features. The PirateNet takes close to three times the time to train at 168 minutes, while the Default takes more than twice the time at 128 minutes, compared to 57 minutes for the Plain configuration. However, it is worth noting that this run time is for the complete 70k training steps. We can see from graph 5 that PirateNet and Default converge much faster at around 25k steps compared to Plain at around 40k.
      Configuration name       $L^{2}_{u}$ error (%)       $L^{2}_{v}$ error (%)       $L^{2}$ norm (%)
      PirateNet       2.10       1.39       2.52
      PirateNet S2S       2.08       1.37       2.49
@@ 4.4 Level set vortex flow
We first tested our three main model configurations, the Plain, the Default, and the PirateNet on the time-reversed vortex flow problem. These configurations represent the use of the same features as in Sections 4.1 to 4.3.2. They were compared using the same basic hyperparameters as seen in Table 11. A particularity with the time-reversed vortex flow is that the flow reverses at the half-period. Therefore, we need to use an S2S scheme with two time windows to ensure the gradients are reset with the proper signs when the flow reverses.
@@ 5 Conclusion
This study has demonstrated the efficacy of advanced PINN architectures such as PirateNet in resolving complex level set problems, particularly those characterized by rigid-body motion and significant interface deformation. Across a series of numerical benchmarks, including Burgers’ equations, Zalesak’s disk, and the time-reversed vortex flow, PirateNet consistently surpassed conventional PINN frameworks in terms of accuracy and computational stability.
@@
These results highlight that the core enhancements intrinsic to PirateNet, such as causal training, sequence-to-sequence learning, random weight factorization, and Fourier feature embeddings, are sufficiently robust for a wide range of level set applications. These enhancements allow PINNs to learn the level set equations without the need for upwind numerical stabilization, geometric reinitialization, or mass conservation schemes as used in classic numerical methods.
In less complex scenarios, such as the 1D and 2D Burgers’ equations, the performance differential between PirateNet and alternative architectures was less pronounced, as all configurations achieved comparably low error rates. However, the utility of PirateNet became unequivocally evident in tackling more sophisticated challenges. For instance, in the time-reversed vortex flow benchmark, PirateNet exhibited a fourfold reduction in error relative to an improved PINN and a tenfold reduction compared to the original PINN. With optimized hyperparameters, the architecture achieved a remarkably low error rate of $L^{2}=0.85\%$.
@@
Future research should prioritize addressing these limitations. Hybrid methodologies that integrate FEM strengths with PINN flexibility or innovations in loss function design explicitly targeting mass conservation could provide promising solutions. Additionally, extending PirateNet’s framework to encompass real-world applications with heightened physical complexity, such as surface tension effects and multifluid interactions, represents an exciting path for future development.
In summary, this study constitutes a substantial advancement in the application of the physics-informed neural networks approach by extending PirateNet’s architecture to accurately model dynamic interfaces without reliance on reinitialization or explicit mass conservation schemes. However, further investigation is imperative to fully bridge the performance gap with traditional numerical solvers and unlock its full potential for both academic inquiry and practical deployment.

### v2

@@ Abstract
This paper advances the use of physics-informed neural networks (PINNs) architectures to address moving interface problems via the level set method. Originally developed for other PDE-based problems, we particularly leverage PirateNet’s features—including causal training, sequence-to-sequence learning, random weight factorization, and Fourier feature embeddings—and tailor them to handle problems with complex interface dynamics. Numerical experiments validate this framework on benchmark problems such as Zalesak’s disk rotation and time-reversed vortex flow. We demonstrate that PINNs can efficiently solve level set problems exhibiting significant interface deformation without the need for upwind numerical stabilization, as generally required by classic discretization methods, or additional mass conservation schemes. However, incorporating an Eikonal regularization term in the loss function with an appropriate weight can further enhance results in specific scenarios. Our results indicate that PINNs with the PirateNet architecture surpass conventional PINNs in accuracy, achieving state-of-the-art error rates of $L^{2}=0.14\%$ for Zalesak’s disk and $L^{2}=0.85\%$ for the time-reversed vortex flow problem, as compared to reference solutions. Additionally, for a complex two-phase flow dam break problem coupling the level set with the Navier-Stokes equations, we propose a geometric reinitialization method embedded within the sequence-to-sequence training scheme to ensure long-term stability and accurate inference of the level set field.
@@ 1 Introduction
This study introduces a PINN-based framework for solving level set equations in moving interface problems and extends it to coupled level set–Navier-Stokes systems with geometric reinitialization for long-term stability. Our contributions include leveraging advanced features of the PirateNet architecture—such as causal training, sequence-to-sequence learning, and random weight factorization—and tailoring them to handle problems with complex interface dynamics.
Our study begins by testing our PINN approach on level set problems, starting with a fabricated simple uniform transport case to highlight the challenges faced by the standard PINN solver and demonstrate the advantages of the proposed techniques. Next, we tackle two well-known level set problems: the rigid-body rotation of Zalesak’s disk [43] and a more complex time-reversed vortex flow problem, which features a varying velocity field [10]. For each case, we compare the reference solutions to results obtained from (a) the original PINN formulation [11], (b) the improved PINN approach [38], and (c) the state-of-the-art PirateNet framework [39]. We finish with a challenging two-phase flow dam break application which couples the level set to the incompressible Navier-Stokes equations. Our findings demonstrate that PirateNet can solve complex level set problems effectively without requiring reinitialization or mass conservation schemes for an exactly conservative flow field. However, for more challenging problems where training can be more unstable, geometric reinitialization can be used to enable long term inference of the level set field.
@@
    With experiments on classical level set benchmarks, we show that PINNs can learn the level set equations without the need for upwind numerical stabilization or mass conservation schemes.
@@
    For complex level set problems were the signed distance property is lost through time, we propose a geometric reinitialization method embedded in the sequence-to-sequence training scheme that allows long term inference of the level set field.
@@ 3.2.1 PirateNet architecture
where each element of the matrix F is sampled from a Gaussian distribution $\mathcal{N}(0,\sigma_{F}^{2})$, and $\sigma_{F}$ is a hyperparameter. Random Fourier features are particularly effective in reducing spectral bias, enabling improved predictions for high-frequency components in the solution [27].
@@ 3.2.2 Random weight factorization
The implementation details of the RWF are as follows [38]. First, the neural network parameters are initialized using a standard scheme, such as the Glorot initialization [54]. Next, for each weight matrix $\bm{W}$, a scale vector $\bm{s}$ is sampled from a multivariate normal distribution $\mathcal{N}(\mu_{rwf},\sigma_{rwf}I)$, with $\mu_{rwf}$ and $\sigma_{rwf}$ are hyperparameters, and $I$ denotes the identity matrix. At initialization, the weight matrix is factorized as $\bm{W}=\text{diag}(\exp(\bm{s}))\cdot\bm{V}$, where $\bm{V}$ represents the remaining factorized parameter. Finally, gradient descent optimization is applied directly to the new parameters, $\bm{s}$ and $\bm{V}$. According to [32], a suitable selection of RWF hyperparameters can shorten the distance between the initialization and the global minimum in the factorized parameter space relative to the original parameter space within their respective loss landscapes. Furthermore, [38] has demonstrated through numerical experiments that, in the case of PINNs, the RWF technique enhances both accuracy and training performance across a range of PDE benchmarks.
@@ 3.2.5 Sequence-to-sequence training
3.2.6 SOAP optimizer
Recently, [56] have shown that traditional first order optimizer like Adam struggle with the composite loss functions of PINNs. The simultaneous optimization of multiple loss terms makes it difficult to align the gradients in the same direction, resulting in competing loss terms and poor optimization. Shampoo with Adam in the Preconditioner’s eigenbasis or SOAP [56, 57] is a second-order quasi-Newton optimizer that naturally aligns gradients to provide better and faster optimization. SOAP is especially effective when solving coupled problems with multiple loss terms.
@@ 4 Results
In the following sections, we present the numerical results obtained from the proposed PINN solver for two well-known level set benchmarks. Before diving into these benchmarks, we validate our PINN framework using two challenging benchmark problems in fluid dynamics, the 1D and 2D Burgers’ equations [61]. To facilitate reading, this section, including ablation studies, was placed in A. Next, we focus on two level set benchmarks: Zalesak’s disk and the time-reversed vortex flow. We finish with a complex application: coupling the level set equation with the incompressible Navier-Stokes equations for modelling of a 2D dam break problem in two-phase flow.
@@ 4.2 Level set vortex flow
We first tested our three main model configurations, the Plain, the Default, and the PirateNet on the time-reversed vortex flow problem. These configurations represent the use of the same features as in Sections A.1 to 4.1.2. They were compared using the same basic hyperparameters as seen in Table 12. A particularity with the time-reversed vortex flow is that the flow reverses at the half-period. Therefore, we need to use an S2S scheme with two time windows to ensure the gradients are reset with the proper signs when the flow reverses.
@@
4.3 Coupled level set-Navier-Stokes equations: dam break flow application
Finally, we consider a very challenging simulation of a 2D dam break problem involving two immiscible, incompressible fluids using a PINN. The flow is governed by the incompressible Navier-Stokes equations coupled with a level set transport equation. The goal is to learn the temporal evolution of the level set, velocity and pressure fields and the interface between both fluids. The computational domain is fully enclosed (four-wall boundary condition), and the only body force acting is gravity.
Let $\Omega\subset\mathbb{R}^{2}$ be a bounded domain with coordinates $(x,y)$ and $t\in[0,T]$ the temporal domain. We define:
  • 1.
    $u(t,x,y),v(t,x,y)$: velocity components in the $x$ and $y$ directions
  • 2.
    $p(t,x,y)$: pressure field
  • 3.
    $\phi(t,x,y)$: level set function describing the interface between fluids
  • 4.
    $\rho(\phi),\mu(\phi)$: fluid density and viscosity defined as functions of $\phi$
  • 5.
    $\mathbf{g}=(0,-g)$: gravitational acceleration acting only in the vertical direction
The level set function is chosen such that:
  • 1.
    $\phi<0$ in fluid 1 (dense fluid)
  • 2.
    $\phi>0$ in fluid 2 (light fluid)
  • 3.
    $\phi=0$ at the interface
Careful consideration needs to be taken to enforce the material properties of both fluids. To differentiate both fluids, we employ a smoothed Heaviside function $H_{\epsilon}(\phi)$:
$H_{\epsilon}(\phi)=\frac{1}{2}\left(1+\tanh\left(\frac{\phi}{\epsilon}\right)\right)$ (47)
This gives smooth transitions for density and viscosity enabling stable gradient flow across the fluid interface:
$\rho(\phi)$ $=\rho_{1}(1-H_{\epsilon}(\phi))+\rho_{2}H_{\epsilon}(\phi)$ (48)
$\mu(\phi)$ $=\mu_{1}(1-H_{\epsilon}(\phi))+\mu_{2}H_{\epsilon}(\phi)$ (49)
4.3.1 Nondimensionalization of Coupled Level Set-Navier-Stokes Equations
In CFD, dimensionless equations are preferred because they simplify equations by removing the dependence on specific units and making them universally applicable at different physical scales. This approach also improves the numerical stability of PINN training by scaling variables to similar orders of magnitude. Using dimensionless equations offers similar benefits as data scaling, a crucial step in traditional machine learning. The dimensionless version of the coupled level set-Navier-Stokes problem is hence used. Furthermore, curriculum training [37] could then be easily used by gradually varying the nondimensional parameters governing the problem.
We define the following characteristic scales for the dam break problem:
  • 1.
    Characteristic Height: $H$ = initial column height of fluid 1
  • 2.
    Characteristic Velocity: $U=\sqrt{g\,H}$ (free fall velocity)
  • 3.
    Characteristic Time: $T=\sqrt{H/g}$ (free fall time)
  • 4.
    Characteristic Pressure: $P=\rho_{\text{ref}}\,g\,H$ (hydrostatic pressure)
The density and viscosity of fluid 1 are used as reference values for the nondimensionalization. The following nondimensional variables are set:
Spatial and Temporal Variables:
$x^{*}=\frac{x}{H},\quad y^{*}=\frac{y}{H},\quad t^{*}=\frac{t}{T}=t\sqrt{\frac{g}{H}}$ (50)
Field Variables:
$u^{*}=\frac{u}{U}=\frac{u}{\sqrt{gH}},\quad v^{*}=\frac{v}{U}=\frac{v}{\sqrt{gH}}$ (51)
$p^{*}=\frac{p}{P}=\frac{p}{\rho_{\text{ref}}gH},\quad\phi^{*}=\frac{\phi}{H}$ (52)
Material Properties:
$\rho^{*}(\phi)=\frac{\rho(\phi)}{\rho_{\text{ref}}},\quad\mu^{*}(\phi)=\frac{\mu(\phi)}{\mu_{\text{ref}}}$ (53)
Using these variables and the chain rule, we can derive the dimensionless equations. Below are the final forms used for training our PINNs.
Non-Dimensional Level Set Residual:
$\mathcal{R}_{\phi}=\frac{\partial\phi^{*}}{\partial t^{*}}+u^{*}\frac{\partial\phi^{*}}{\partial x^{*}}+v^{*}\frac{\partial\phi^{*}}{\partial y^{*}}$ (54)
Non-Dimensional Continuity Residual:
$\mathcal{R}_{\text{cont}}=\frac{\partial u^{*}}{\partial x^{*}}+\frac{\partial v^{*}}{\partial y^{*}}$ (55)
Non-Dimensional X-Momentum Residual:
$\mathcal{R}_{\text{mom-x}}&=\frac{\partial u^{*}}{\partial t^{*}}+u^{*}\frac{\partial u^{*}}{\partial x^{*}}+v^{*}\frac{\partial u^{*}}{\partial y^{*}}+\frac{1}{\rho^{*}(\phi)}\frac{\partial p^{*}}{\partial x^{*}}\\ &-\frac{1}{\text{Re}}\cdot\frac{\mu^{*}(\phi)}{\rho^{*}(\phi)}\left(\frac{\partial^{2}u^{*}}{\partial(x^{*})^{2}}+\frac{\partial^{2}u^{*}}{\partial(y^{*})^{2}}\right)-g^{*}_{x}\end{aligned}$ (56)
Non-Dimensional Y-Momentum Residual:
$\mathcal{R}_{\text{mom-y}}&=\frac{\partial v^{*}}{\partial t^{*}}+u^{*}\frac{\partial v^{*}}{\partial x^{*}}+v^{*}\frac{\partial v^{*}}{\partial y^{*}}+\frac{1}{\rho^{*}(\phi)}\frac{\partial p^{*}}{\partial y^{*}}\\ &-\frac{1}{\text{Re}}\cdot\frac{\mu^{*}(\phi)}{\rho^{*}(\phi)}\left(\frac{\partial^{2}v^{*}}{\partial(x^{*})^{2}}+\frac{\partial^{2}v^{*}}{\partial(y^{*})^{2}}\right)-g^{*}_{y}\end{aligned}$ (57)
We note that to reduce computational complexity and enable more stable training, only the dominant second-order viscous terms are retained for the momentum equations. The terms involving derivation of $\mu$ are dropped under the assumption that the residuals aren’t computed for collocation points exactly on the sharp interface. There is an implicit continuity of the total Cauchy stress across the interface.
4.3.2 Problem setup
The fluid properties for this problem are resumed in Table 5. The Heaviside function 47 is smoothed with $\epsilon=0.01$ while gravity is set as $g=9.81\frac{m}{s^{2}}$, acting in the negative $y$ direction. The spatial domain is defined as a square box of length $L=1.168\,m$ on each side and the simulation is run for $0.8$ seconds. The denser Fluid 1 is initially contained in a rectangle of width $0.292\,m$ and height $0.584\,m$. Hence, the nondimensional constants are set as $H=0.584\,m$, $U=2.4\,\frac{m}{s}$, $T=0.24\,s$ and $P=5729\,Pa$. Using these values and using the fluid properties of the denser fluid, the reference Reynolds number is computed as $Re=\frac{\rho_{1}*\sqrt{g*H}*H}{mu_{1}}=35*10^{3}$.
      Fluid       Density $\rho$ ($\frac{kg}{m^{3}}$)       Dynamic viscosity $\mu$ ($Pa*s$)
      Fluid 1 (dense)       1000       0.04
      Fluid 2 (light)       50       0.02
Table 5: Coupled level set-NS: fluid properties.
The reference solution was generated using the same FEM solver as previous sections. The results were interpolated on a $41\times 41$ grid for $160$ time steps.
4.3.3 Geometric reinitialization
As with classical level set solvers based on the FEM, the signed distance property of the level set function is gradually lost during the training of our PINN. Although regularization with an Eikonal loss term helps enforce the signed distance property to some extent, we observed that it was insufficient to maintain stability over long time horizons.
To address this, we propose a simple geometric reinitialization step that is applied between time windows of sequence-to-sequence training. Before training begins on the new window of sequence-to-sequence, we reinitialize the level set field to restore its signed distance property. The geometric reinitialization method is based on the Euclidean distance transform (EDT), a function already available in Scipy [63]. This method produces a new level set field where the zero level set remains in place, but the values on either side are reset to their signed distances from the interface.
This approach was found to significantly improve the stability and convergence of training for the coupled Navier-Stokes and level set PINN solver, especially when dealing with high density ratios and sharp interface dynamics.
4.3.4 Results
Given the computational cost of this problem, a Modified MLP was used instead of the PirateNet. Generally the cost of PirateNet is about three times that of the Modified MLP. This allows to fit more collocation points in the memory of a single V100 GPU. The SOAP optimizer is used due to the complex nature of this coupled problem with several conflicting loss terms, three inputs and four outputs. A more detailed list of hyperparameters can be found in Table 13. With the proposed method and hyperparameters, the PINN obtained a relative $L^{2}$ error when compared to the FEM reference of $L^{2}=5.6\%$. We note that the frames considered in this error calculation are only the ones subjected to the reinitialization step. Figure 12 shows the evolution of the reference and predicted level set fields.
Figure 12: Coupled level set-NS: Temporal evolution of $\phi$ for the FEM reference and PINN
It can be seen that the PINN solver generally captures the evolution of the level set field but struggles with the sharper details such as splashes. This error could probably be lowered with more collocation points and time windows, but this would significantly increase computational cost.
Furthermore, the mean absolute error of mass loss is $Mass_{MAPE}=3.8\%$. The evolution of mass loss plotted in Figure 13 shows that the PINN suffers from large mass loss after the sudden drop of the dense fluid column, but is capable of restoring it’s mass in time. A significant pattern of error accumulation isn’t noticed.
Figure 13: Coupled level set-NS: Temporal evolution of mass loss
A significant downfall of solving this problem with a PINN is the computational requirements. The PINN took 16 hours to train on a single V100 GPU while the reference FEM took 1.25 hours on a laptop CPU.
Nevertheless, this example shows that PINNs can capture the evolution of a level set field for a complex coupled level set-Navier-Stokes problem.
@@ 5 Conclusion
This study has demonstrated the efficacy of advanced PINN architectures such as PirateNet and appropriate training strategies in resolving complex level set problems, particularly those characterized by rigid-body motion and significant interface deformation. Across a series of numerical benchmarks, including Burgers’ equations, Zalesak’s disk, and the time-reversed vortex flow, PirateNet consistently surpassed conventional PINN frameworks in terms of accuracy and computational stability. A PINN with the Modified MLP architecture also successfully solved a complex dam break problem in two phase flow by coupling the level set equation with the Navier-Stokes equations.
@@
These results highlight that the core enhancements intrinsic to PirateNet, such as causal training, sequence-to-sequence learning, random weight factorization, and Fourier feature embeddings, are sufficiently robust for a wide range of level set applications. These enhancements allow PINNs to learn the level set equations without the need for upwind numerical stabilization or mass conservation schemes as used in classic numerical methods. However, in the case of more complex problems like the tested dam break where the signed distance property needs to be reinitialized, a geometric reinitialization can be introduced between time windows of sequence-to-sequence training to allow long time inference.
In less complex scenarios, such as the 1D and 2D Burgers’ equations, the performance differential between PirateNet and alternative architectures was less pronounced, as all configurations achieved comparably low error rates. However, the utility of PirateNet became unequivocally evident in tackling more sophisticated challenges. For instance, in the time-reversed vortex flow benchmark, PirateNet exhibited a fourfold reduction in error relative to an improved PINN and a tenfold reduction compared to the original PINN. With optimized hyperparameters, the architecture achieved a remarkably low error rate of $L^{2}=0.85\%$. However, PirateNet requires more memory than alternatives like the Modified MLP which makes PirateNet harder to use in limited memory application.
@@
Future research should prioritize addressing these limitations. Hybrid methodologies that integrate FEM strengths with PINN flexibility or innovations in loss function design explicitly targeting mass conservation could provide promising solutions. Additionally, extending PirateNet’s framework to encompass real-world applications with heightened physical complexity represents an exciting path for future development.
In summary, this study constitutes a substantial advancement in the application of the physics-informed neural networks approach by extending PirateNet’s architecture to accurately model dynamic interfaces without reliance on upwind numerical stabilization or explicit mass conservation schemes. However, further investigation is imperative to fully bridge the performance gap with traditional numerical solvers and unlock its full potential for both academic inquiry and practical deployment.